OpenStudy (narissa):

x+4=12 A -16 B 16 C 8 D -8 -7=21+X A 28 B 14 C -14 D -28 7x=-21 A-28 B-21 C-3 D-14

OpenStudy (narissa):

@MyNameIsNemo

OpenStudy (mynameisnemo):

On question 1. you want to isolate the variable (get x alone), to do so, you need to move the 4 to the other side by subtracting 4 from both sides x + 4 = 12 - 4 - 4 x= 12-4 x=8

OpenStudy (mynameisnemo):

do the same for question 2; the equation may be backwards, but same basic steps. -7 = 21 + X - 21 - 21 -7 - 21 = X -28 = X

OpenStudy (mynameisnemo):

Now, you still have the same reasoning for each step on question 3 but because X is being multiplied by 7, you need to do the opposite, which would be to divide it. 7x = -21 /7 /7 X = -3

OpenStudy (narissa):

thanks a bunch i will tag you in my next post